• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Solvignon

    Log entry time 16:05:58 on March 7,2009

    Entry number 265146

    keyword=Shift summary


    DAY shift summary: P. Solvignon (SL), A. Deur (TO), Nuruzzaman (3rd)
    =========================================================================

    EVENTS
    ------
    08:00 took over just after BHWP changed from OUT to IN
    11:50 NMR measurement: 57.32%
    14:40 MCC called about their problem with beam delivery. They are waiting for
    the expert to come in and change a power supply card. In the meantime they can send us beam while waiting for the expert.
    15:54 RC called to ask that we lower the beam current because the deadtime is too high.
    15:58 MCC called to start the card change in a power supply.

    RUNS
    ----

    Bigbite:

    08:01-08:50 #2040 prod, good
    08:50-09:43 #2041 prod, good
    09:43-10:43 #2042 prod, good
    10:44-11:30 #2043 prod, good
    11:31-11:49 #2044 prod, good
    11:51-12:39 #2045 prod, good
    12:40-12:49 #2046 prod, good --> stopped because of beam problem
    12:56-12:58 #2047 prod, good --> stopped because of beam problem
    13:11-14:06 #2048 prod, good
    14:06-14:26 #2049 prod, good --> stopped because of beam problem
    14:37-15:32 #2050 prod, good
    15:32-15:56 #2051 prod, good
    15:59- #2052 prod

    HRSL:

    08:01-11:48 #20578 prod, good
    11:51-12:49 #20579 prod, good
    12:56-12:58 #20580 prod, good --> stopped because of beam problem
    13:11-14:26 #20581 prod, good --> stopped because of beam problem
    14:37-15:56 #20582 prod, good --> stopped because of beam problem
    15:59- #20583